Operations Report – April 10th

Crew 314 Operations Report 10Apr2025

SOL: 4

Name of person filing report: Antoine (Crew Engineer)

Non-nominal systems :

The toilet repair performed yesterday is still holding and functioning properly.

The CO detector in the science dome is not working.

Notes on non-nominal systems : /

ROVERS

Spirit rover used: /

Hours: /

Beginning charge: /

Ending charge: /

Currently charging: /

Opportunity rover used:

Hours: 214.8 (0.4 used)

Beginning charge: 100%

Ending charge: 65%

Currently charging: No

Curiosity rover used:

Hours: 306.9 (0,4 used)

Beginning charge: 100%

Ending charge: 73%

Currently charging: No

Perseverance rover used:

Hours: /

Beginning charge: /

Ending charge: /

Currently charging: /

General notes on rovers: Spirit is still out of use due to a battery issue.

Summary of Hab operations : Today in the Habitat, we focused on cooking, socializing, completing our individual reports, and working on our respective experiments. Additionally, we did a bit of physical exercise to stay in shape. Finally, we held our usual team meeting to discuss progress and upcoming tasks.

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 7 p.m.
1) Per formula (see notes): 15.55 gallons
2) Smart Home Dashboard (see notes): 15.08 gallons

Water (static tank, remaining gallons): 374.99 gallons (per formula) & 392.88 gallons (Smart Home Dashboard)

Static tank pipe heater (on or off): on

Static tank heater (on or off) : on

Toilet tank emptied (no or yes): no

Summary of internet: used for music playing, mails, research for experiments, learning languages.

Summary of suits and radios (EVA #6) :

Suits : 2, 4 and 5.

Radio : 3, 4 and 5.

Summary of GreenHab operations: put notes here: I watered all the plants with water.

The door of the GreenHab was left open for 3 hours 35 minutes (2.10 pm to 5.45 pm) to reduce the temperature inside (maximum of 98.8 °F)

WATER USE: 9 gallons

Heater (On or Off): Heater on (setup temperature = 72°F)

Supplemental light (hours of operation): 7-10 pm

Harvest (name, weight in grams): 4 g of parsley and 6 g of basilic

Summary of ScienceDome operations: put notes here

Inoculation of Esherichia coli on 96 well plate + UV-C exposure

Inoculation of Bacillus thuringiensis on petri dishes + soil+ UV-C exposure

Isolation of arbuscular fungi spores (after exposure to stresses)

Disinfection of arbuscular fungi spores

Spores on a membrane to test the germination after exposure to stresses

Plantain seedlings (use for experiment) watering

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: (Tools used, work done)

At 9:20 a.m., the Crew Engineer took 2 zip ties and a dark ribbon to repair equipment in the field if needed. However, no repair was necessary, and all materials were returned to their place.

Around noon, the HSO Officer went to the RAM to take 2 inches of blue tape to label a container and indicate its intended contents.

Around 4:00 p.m., the Crew Engineer took 6 small zip ties, 3 large ones, 50 inches of wire, and 20 inches of grey vinyl duct tape to reattach tarps and seal holes.
Summary of any observatory issues: None for the moment.

Summary of health and safety issues: Everyone was in good health today, with no issues reported. There were no safety concerns to note.

Questions, concerns and requests to Mission Support: None for the moment.

Operations Report – April 11th

Crew 314 Operations Report 11-04-2025

SOL: 5

Name of person filing report: Antoine Dubois (Crew Engineer)

Non-nominal systems :

The CO detector is working.

Notes on non-nominal systems :

CO detector: The battery has been replaced.

ROVERS

Spirit rover used:

Hours: (before EVA) /

Beginning charge: (Before EVA) /

Ending charge: (On return from EVA, before recharging) /

Currently charging: yes

Opportunity rover used:

Hours: 215.7 (0.9 hours used)

Beginning charge: 100%

Ending charge: 47%

Currently charging: no

Curiosity rover used:

Hours: 307.6 (0.7 hours used)

Beginning charge: 100%

Ending charge: 67%

Currently charging: no

Perseverance rover used:

Hours: /

Beginning charge: /

Ending charge: /

Currently charging: yes

General notes on rovers:

Spirit can be used but for short EVAs, perseverance is usable as well.

We are concerned about Opportunity’s battery ? Because it’s the second time that the battery percentage goes down really fast.

Summary of Hab operations: Today in the Habitat, we focused on cooking, socializing, completing our individual reports, and working on our respective experiments. Additionally, we did a bit of physical exercise to stay in shape. We also conducted the third medical emergency simulation (Odile’s experiment), and it went really well. Finally, we held our usual team meeting to discuss progress and upcoming tasks.

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 7 p.m.
1) Per formula (see notes): 17.78 gallons
2) Smart Home Dashboard (see notes): 19.90 gallons

Water (static tank, remaining gallons): 357.21 gallons (per formula) & 372.98 gallons (Smart Home Dashboard)

Static tank pipe heater (on or off): off

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): yes

Summary of internet: used for music playing, mails, research for experiments, learning languages.

Summary of suits and radios:

suits : 10, 9, 8, 7

radios : 4, 1, 6, 7, 3, 2

Summary of GreenHab operations: put notes here:

I watered all the plants with water.

The door of the GreenHab was left open for 4 hours 25 minutes (1.50 pm to 6.15 pm) to reduce the temperature inside (maximum of 100.9 °F).

I sowed the seeds of sprouts of China Rose radish, Kale, Arugula and cress seeds in 3 small rectangular black pots.

I took 3 garden labels for the new cultures.

WATER USE: 11 gallons

Heater (On or Off): Heater on (setup temperature = 72°F)

Supplemental light (hours of operation): 7-10 pm

Harvest (name, weight in grams): 1 g of basilic

Summary of ScienceDome operations: put notes here

Check the condition of plantain seedlings and watering

Check Petri dishes for arbuscular fungi spore germination

Milieu preparation

Exposure to UV-C of Bacillus thuregiensis

Dual split (Heat or AC, On or Off): OFF

Summary of RAM operations: (Tools used, work done)

At 9:15 a.m., the Crew Engineer took 2 zip ties and a dark ribbon to repair equipment in the field if needed. However, no repair was necessary, and all materials were returned to their place.

Batoul picked up the package that was delivered for her experiment.

Summary of any observatory issues: Nothing regarding the hardware. For the software part, there is still work to be done in order to master every aspect of it (please refer to the dedicated report for further details).

Summary of health and safety issues: Everyone was in good health today, with no issues reported. There were no safety concerns to note.

Questions, concerns and requests to Mission Support: None for the moment.

Operations Report – April 12th

Crew 314 Operations Report 12-04-2025

SOL: 6

Name of person filing report: Antoine Dubois (Engineer)

Non-nominal systems :

The tunnels have been checked.

Notes on non-nominal systems :

Tunnels: We carried out a complete check of the tunnels due to the wind speed and 11 cable ties have been replaced. The 100 pack of cable ties (11 inch) is empty.

ROVERS

Spirit rover used: /

Hours: (before EVA)/

Beginning charge: (Before EVA) /

Ending charge: (On return from EVA, before recharging) /

Currently charging: yes

Opportunity rover used: Not used

Hours: /

Beginning charge: /

Ending charge: /

Currently charging: yes

Curiosity rover used: EVA#8

Hours: 307.6

Beginning charge:100

Ending charge: 59

Currently charging: No

Perseverance rover used: EVA#8

Hours: 318.7

Beginning charge: 100

Ending charge: 58

Currently charging: No

General notes on rovers: Has Opportunity been checked out regarding the battery ? It seemed to lower quickly during our previous EVAs as mentioned yesterday in our report.

Summary of Hab operations : Today in the Hab, we dedicated time to cooking, socializing, making progress on our experiments and advancing our individual reports. We also did some physical exercise to maintain our shape. In addition, we filmed a few videos for our social media channels. To wrap up the day, we held our regular team meeting to review our progress and plan the tasks ahead.

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 7 p.m.
1) Per formula (see notes): 13.20 gallons
2) Smart Home Dashboard (see notes): 17.30 gallons

Water (static tank, remaining gallons): 344.02 gallons (per formula) & 355.68 gallons (Smart Home Dashboard)

Static tank pipe heater (on or off): off

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): no

Summary of internet: Research for experiments, languages learning and music listening

Summary of suits and radios:

Suits EVA#8 : 7, 8, 9 and 10

Radios EVA #8 : 3, 4, 6 and 7

Summary of GreenHab operations: put notes here:

I watered all the plants with water in which I added 2 teaspoons of Miracle-Gro (I finished one small bag) per 2 gallons of water.

The door of the GreenHab was left open for 7 hours 50 minutes (11.10 am to 7 pm) to reduce the temperature inside (maximum of 102.7 °F).

I watered and checked the seeds of the 3 small rectangular black pots (sprouts of China Rose radish) with the new cultures (Kale, Arugula and cress). The seeds of sprouts of China Rose radish (which were in water to initiate germination) were sown in a small rectangular black pot and I added a label.

I removed plants in the square blue pot because they were at the end of their life.

I cleaned the floor of the GreenHab.

WATER USE: 8.75 gallons

Heater (On or Off): Heater on (setup temperature = 72°F)

Supplemental light (hours of operation): 7-10 pm

Harvest (name, weight in grams): 1 rainbow cherries and 1 spring onion

Summary of ScienceDome operations: put notes here

Check the condition of plantain seedlings and watering

Check Petri dishes for arbuscular fungi spore germination

Shooting videos explaining our experiments for our social networks

Isolation of AMF spores

Inoculation of Esherichia coli

Inoculation on petri dishes of Bacillus Thuringiensis with soils

Treatment of treated dishes

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations:

At 9:20 a.m., the Crew Engineer took 2 zip ties and a dark ribbon to repair equipment in the field if needed. However, no repair was necessary, and all materials were returned to their place.

We shot some videos explaining our experiments for our belgian social network during the afternoon.

At 7:10 p.m., we took 11 cable ties in the RAM to fix the tunnel.

Summary of any observatory issues: None, weather conditions were not good enough, too many clouds and strong winds making it unsafe.

Summary of health and safety issues: Everyone was in good health today, with no issues reported. There were no safety concerns to note.

Questions, concerns and requests to Mission Support:

A request was submitted to Mission Support to assist the crew in closing the main entrance door, which had suddenly slammed open due to strong winds and became stuck. Thanks to Mission Support’s prompt intervention, the issue was successfully resolved.

Can we have a new pack of 11 inch cable ties ? The package was emptied during the tunnel check.

Since some of the canned goods and food supplies (such as eggs, coffee, etc.) were already partially used upon our arrival, would it be possible to receive a resupply?

Supplemental Operations Report – April 2025

Date: 4/12/2025
Name of person filing report: Sergii Iakymov
Reason for Report: Routine
Non-Nominal Systems: Crew car. Power system battery, inverters, generator. Robotic observatory. HAB outer shell. Spirit.

Power system:
Solar: The battery bank does not hold charge when sun is down and low on the horizon. Inverter Slave 1 and 2 went offline and do not restart, which limits us to 5kW when on solar.
Main generator has been monitored for oil leaks; minor leaks observed, Moreover, generator is consuming oil because of worn piston rings. Adding oil every day is necessary. Generator is limited to 8kW, see previous reports for details. Engine governor is being stuck in open position during startups and needs to be replaced, new part ordered.
Main generator:
1) Oil, oil filter and air filter changed on 04/11/2025. Hours: 9260.2
2) Current hours – 9274.4.2

Propane Readings:
Station Tank: 74%
Director Tank: 69%
Intern Tank: 88%
Generator Tank: 64%

Water:
Hab Static Tank – 344 gallons
GreenHab – 144 gallons
Outpost tank – 400 gallons

Rovers:
Sojourner rover used: Yes
Hours: 212.7
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: No
Notes on Rovers: Spirit is not charging. Perseverance’s parking brake needs to be adjusted.

Cars:
Hab Car used and why, where: To Hanksville for supplies.
Crew Car used and why, where: By crew from Grand Junction.
General notes and comments: Crew car driver’s side front ball joint is bad. Low oil pressure is most likely caused by a bad sensor (it does not change due to RPM or temperature like it should).

Summary of Internet: Nominal
EVA suits and radios:
Suits: Nominal.
Comms: Nominal.
T-Echo EVA-link: Nominal

Campus wide inspection, if action taken, what and why: Nominal
Summary of Hab Operations: Small cracks in the wall of the Hab in the loft area are allowing cold air intrusion. Static tank and pipe heaters unplugged. Thermostat setup for 60F due to warm weather.
Summary of GreenHab Operations: Nominal
Summary of SciDome Operations: Dual split is being turned on during day time for cooling.
Summary of Observatories Operations: Robotic observatory offline.
Summary of RAM Operations: Nominal
Summary of Outpost Operations: Static tank heater unplugged and removed. Screen doors installed in both trailers. Intern trailer heater turned off due to warm weather.
Summary of Health and Safety Issues: Nominal

Operations Report – April 13th

Crew 314 Operations Report 13-04-2025

SOL: 7

Name of person filing report: Antoine Dubois (Engineer)

Non-nominal systems :

The number of cable ties have been checked.

Notes on non-nominal systems :

Cable ties: There are 4.5 packs of 50 zip ties (24"), approximately 50 zip ties (8"), and around 15 zip ties (6") in storage.

ROVERS

Spirit rover used : /

Opportunity rover used: EVA # 9

Hours: 215.8 (0.8 used).

Beginning charge: 100%

Ending charge: 44%

Currently charging: No

Curiosity rover used: EVA # 9

Hours: 308.3 (0.8 used)

Beginning charge: 100%

Ending charge: 65%

Currently charging: No

Perseverance rover used: /

General notes on rovers: /

Summary of Hab operations: Today in the Hab, we dedicated time to cooking, socializing, making progress on our experiments and advancing our individual reports. We also did some physical exercise to maintain our shape. During the afternoon, the fourth medical simulation took place, included in Odile’s experiment. To wrap up the day, we held our regular team meeting to review our progress and plan the tasks ahead.

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 7 p.m.
1) Per formula (see notes): 17.90 gallons
2) Smart Home Dashboard (see notes): 16.56 gallons

Water (static tank, remaining gallons): 326.11 gallons (per formula) & 339.12 gallons (Smart Home Dashboard)

Static tank pipe heater (on or off): off

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): no

Summary of internet: Research for experiments, languages learning and music listening

Summary of suits and radios:

Radio : 3, 4, 5 et 6.

Suits : 7, 8, 9 et 10.

Summary of GreenHab operations: put notes here

I watered all the plants with water.

The door of the GreenHab was left open for 8 hours and 15 minutes (9.10 am to 5.25 pm) to reduce the temperature inside (maximum of 108.3 °F).

I watered and checked the seeds of the 3 small rectangular black pots (sprouts of China Rose radish) with the new cultures (Kale, Arugula and cress).

I removed Arugula plants because the plants were too old to be eaten.

I cut the dead branches of the tomato plants and tied up the overhanging branches.

I added stakes with a sticky paper to try to catch the drosophila present in the GreenHab.

I tied up the dill branches with a rope and tried to separate the different cultures (basil, dill, parsley, spring onions, chives, garlic) in the big blue pot.

I cleaned the floor of the GreenHab.

WATER USE: 9.5 gallons

Heater (On or Off): Heater on (setup temperature = 72 °F)

Supplemental light (hours of operation): 7-10 pm

Harvest (name, weight in grams): 6 spring onions, 8 rainbow cherries and 6 g of parsley

Summary of ScienceDome operations: put notes here

Check the condition of plantain seedlings and watering

Inoculation of soil samples (from the EVA) with spores of arbuscular fungi

Application of different stress (UV, gravity, temperature and control) conditions to soil samples (from EVA)

Batoul still made notable progress, adding a colored marker to each of her samples for future identification and analysis

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: We picked up the food cans and counted the number of cable ties at 5 p.m.

Summary of any observatory issues: The solar observatory was used in the morning. In the afternoon, cloud cover prevented further observations.

Summary of health and safety issues: Everyone was in good health today, with no issues reported. There were no safety concerns to note.

Questions, concerns and requests to Mission Support: Food requested by the crew was provided by the mission support

Operations Report – April 14th

Crew 314 Operations Report 14-04-2025

SOL: #8

Name of person filing report: Antoine Dubois (Engineer)

Non-nominal systems :

Check that tunnels between modules are in good condition.

Problem with Curiosity rover.

Notes on non-nominal systems :

All tunnels are in good condition.

Curiosity seems to have a steering problem. Actually the steering alignment is off and if you don’t hold the wheel, the rover drifts to the left.

ROVERS

Spirit rover used:

Hours: 273.2 (0.2 used)

Beginning charge: 100

Ending charge: 90

Currently charging: no

Opportunity rover used:

Hours: 216.6 (0,3 used)

Beginning charge: 100%

Ending charge: 75%

Currently charging: No

Curiosity rover used:

Hours: 309.1 (0.2 used)

Beginning charge: 100

Ending charge: 91

Currently charging: no

Perseverance rover used:

Hours: 319.6 (0,4 used)

Beginning charge: 100%

Ending charge: 87%

Currently charging : No

General notes on rovers:

Summary of Hab operations: Today in the Hab, we dedicated time to cooking, socializing, making progress on our experiments and advancing our individual reports. We also did some physical exercise to maintain our shape. To wrap up the day, we held our regular team meeting to review our progress and plan the tasks ahead.

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 6pm
1) Per formula (see notes): 23.13 gallons
2) Smart Home Dashboard (see notes): 19.50 gallons

Water (static tank, remaining gallons): 302.98 gallons (per formula) & 319.62 gallons (Smart Home Dashboard)

Static tank pipe heater (on or off): off

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): yes

Summary of internet: Research for experiments, languages learning and music listening.

Summary of suits and radios :

EVA #10 :

Suits : 7, 8, 9 and 10

Radios : 1, 2, 5 and 7.

EVA #11 :

Suits : 7, 8, 9 and 10.

Radios : 3, 4, 5 and 6.

Summary of GreenHab operations: put notes here:

I watered all the plants with water.

The door of the GreenHab was left open for 7 hours and 55 minutes (9.30 am to 5.25 pm) to reduce the temperature inside (maximum of 95 °F).

WATER USE: 10.5 gallons

Heater (On or Off): Heater on (setup temperature = 72 °F)

Supplemental light (hours of operation): 7-10 pm

Harvest (name, weight in grams): None

Summary of ScienceDome operations:

Check the condition of plantain seedlings and watering

Check the presence or absence of arbuscular fungi spores in the soil sampled during the EVAs

Counting CFU in petri dishes of Bacillus thuringiensis group 1 and 2

Exposure to UV-C of Bacillus thuringiensis group 3

Exposure to UV-C of Esherichia coli group 3

Staining of biofilms with cristal violet

Observing Bacillus thuringiensis with the microscope

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations:

At 1:30 p.m., the Crew Engineer took 2 zip ties and a dark ribbon to repair equipment in the field if needed. However, no repair was necessary, and all materials were returned to their place.

Summary of any observatory issues: The astronomer participated in both EVAs today for a personal experiment, so no astronomical observations were conducted.

Summary of health and safety issues: Cf HSO report. There were no safety concerns to note.

Questions, concerns and requests to Mission Support: Crew was wondering if Mission Support would be interested in having a stationary bike at the station for future crews, particularly for potential physical activity research. The bike can be disassembled and stored in its original box, along with the necessary tools for easy reassembly.

Operations Report – April 12th

Crew 314 Operations Report 15-04-2025

SOL: #8

Name of person filing report: Antoine Dubois (Engineer)

Non-nominal systems :

Arnaud repaired a watering can.

Notes on non-nominal systems :

The watering can was worn and repaired with tape.

ROVERS

Spirit rover used:

Hours: /

Beginning charge: /

Ending charge: /

Currently charging: yes

Opportunity rover used: /

Hours: /

Beginning charge: /

Ending charge: /

Currently charging: yes

Curiosity rover used:

Hours: 309.3 (used 0.9)

Beginning charge: 100%

Ending charge: 56%

Currently charging: no

Perseverance rover used:

Hours: 320.0 (used 0.9)

Beginning charge: 100%

Ending charge: 60%

Currently charging: no

General notes on rovers: /

Summary of Hab operations: Today in the Hab, we began our day with the fifth medical simulation, which was a bit different from the previous ones as it took place at the end of the night to introduce stress-inducing conditions. Afterward, we dedicated time to cooking, socializing, making progress on our experiments, and advancing our individual reports. We also engaged in some physical exercise to maintain our fitness. To wrap up the day, we held our regular team meeting to review our progress and plan the tasks ahead.

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 6 p.m.
1) Per formula (see notes): 20.51 gallons
2) Smart Home Dashboard (see notes): 20.92 gallons

Water (static tank, remaining gallons): 282.46 gallons (per formula) & 298.70 gallons (Smart Home Dashboard)

Static tank pipe heater (on or off): off

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): no

Summary of internet: Research for experiments, languages learning and music listening.

Summary of suits and radios:

Suits : 7, 8, 9 and 10.

Radios : 3, 4, 5 and 7.

Summary of GreenHab operations: put notes here

I watered all the plants with water.

The door of the GreenHab was left open for 4 hours and 20 minutes (2.00 am to 6.20 pm) to reduce the temperature inside (maximum of 99.5 °F).

Repair of pink sprinkler (broken top).

WATER USE: 10.75 gallons

Heater (On or Off): Heater on (setup temperature = 72 °F)

Supplemental light (hours of operation): 7-10 pm

Harvest (name, weight in grams): 126 g of basil

Summary of ScienceDome operations: put notes here

Treatment of Bacillus thuringiensis

Watering plantain plantlets

Dual split (Heat or AC, On or Off): AC, ON

Summary of RAM operations: /

Summary of any observatory issues: No observation today.

Summary of health and safety issues: Everyone was in good health today, with no issues reported. There were no safety concerns to note.

Questions, concerns and requests to Mission Support: /

Operations Report – April 16th

Crew 314 Operations Report 16-04-2025

SOL: 10

Name of person filing report: Antoine Dubois (Engineer)

Non-nominal systems : /

Notes on non-nominal systems : /

ROVERS

Spirit rover used:

Hours : 273.4 (0.5 hours used)

Beginning charge : 90%

Ending charge: 67%

Currently charging: no

Opportunity rover used:

Hours: 216.9 (0.7 hours used)

Beginning charge: 100%

Ending charge: 61%

Currently charging: no

Curiosity rover used:

Hours: /

Beginning charge: /

Ending charge: /

Currently charging: yes

Perseverance rover used:

Hours: /

Beginning charge: /

Ending charge: /

Currently charging: /

General notes on rovers : yes

Summary of Hab operations: Today in the Hab, we dedicated time to cooking, socializing, making progress on our experiments, and advancing our individual reports. We also engaged in some physical exercise to maintain our fitness. To wrap up the day, we held our regular team meeting to review our progress and plan the tasks ahead.

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 6 pm
1) Per formula (see notes): 16.08 gallons
2) Smart Home Dashboard (see notes): 18.97 gallons

Water (static tank, remaining gallons): 266.38 gallons (per formula) & 279.73 gallons (Smart Home Dashboard)

Static tank pipe heater (on or off): off

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): no

Summary of internet: Research for experiments, mail, languages learning and music listening.

Summary of suits and radios :

Suits : EVA 13 : 7, 8, 9, 10 & EVA 14 : 10, 9, 8

radios : EVA 13 : 1, 5, 6, 7 & EVA 14 : 3,5,4

Summary of GreenHab operations: put notes here

I watered all the plants with water.

The door of the GreenHab was left open for 3 hours and 50 minutes (1:00 am to 4:50 pm) to reduce the temperature inside (maximum of 97.3 °F).

WATER USE: 8 gallons

Heater (On or Off): Heater on (setup temperature = 72 °F)

Supplemental light (hours of operation): 7-10 pm

Harvest (name, weight in grams): 6 g of parsley, 59 g of rainbow cherries

Summary of ScienceDome operations: put notes here

Observing the results in the microscope

End of experiment with UV and gravity (dismantling of equipment)

Isolation of membranes and staining of the spores to observe germination

Dual split (Heat or AC, On or Off): AC, ON

Summary of RAM operations: No operations were carried out in RAM.

Summary of any observatory issues: No observation today.

Summary of health and safety issues: Everyone was in good health today, with no issues reported. There were no safety concerns to note.

Questions, concerns and requests to Mission Support: None

Operations Report – April 17th

Crew 314 Operations Report 17-04-2025

SOL: 11

Name of person filing report: Antoine Dubois (Engineer)

Non-nominal systems :

Tunnel inspection and some damage due to wind.

Suit #3 has been repaired.

Notes on non-nominal systems :

A tunnel inspection was conducted in the early afternoon, and all coverings were found to be in good condition. A second check was carried out at 6:30 PM, during which one tarp near the RAM was found to be damaged. It was reattached, but should be monitored closely in the coming days due to ongoing wind conditions. Approximately 15 zip ties were used to secure other sections of the tunnel, and no additional damage was observed.

Suit #3 : The missing part was hidden at the top of the shelf. The part has been attached to the suit. The suit is operational.

ROVERS

Spirit rover used: /

Opportunity rover used: /

Curiosity rover used:

Hours: 310.2 (0,9 used)

Beginning charge: 100%

Ending charge: 50 %

Currently charging:No

Perseverance rover used:

Hours: 320.9 (1 used)

Beginning charge: 100%

Ending charge: 49%

Currently charging: No

General notes on rovers: /

Summary of Hab operations: Over the past 24 hours, we participated in our final medical emergency simulation. The scenario began at midnight, when we were woken by an alarm in the Hab and directed to the GreenHab, where the simulation took place. After completing the exercise, we returned to the Hab for a debriefing session and were back in bed by 1:00 AM.

Later in the day, we spent time cooking, socializing, advancing our experiments, and working on individual reports. We also engaged in physical exercise to maintain our fitness. To conclude the day, we held our regular team meeting to review our progress and plan the tasks ahead.

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 7 pm
1) Per formula (see notes): 15.03 gallons
2) Smart Home Dashboard (see notes): 19.31 gallons

Water (static tank, remaining gallons): 251.35 gallons (per formula) & 260.42 gallons (Smart Home Dashboard)

Static tank pipe heater (on or off): off

Static tank heater (on or off) : off

Toilet tank emptied (no or yes): yes

Summary of internet: Research for experiments, mail, languages learning and music listening.

Summary of suits and radios:

Suits : 7, 8, 9 and 10.

Radio : 3, 4, 5 and 6.

Summary of GreenHab operations:

I watered all the plants with water.

The door of the GreenHab was left open for 4 hours and 30 minutes (9:10 am to 1:40 pm) to reduce the temperature inside (maximum of 94.6 °F).

WATER USE: 9 gallons

Heater (On or Off): Heater on (setup temperature = 72 °F)

Supplemental light (hours of operation): 7-10 pm

Harvest (name, weight in grams):12g of mint, 15 g of parsley

Summary of ScienceDome operations: put notes here:

Staining biofilms with crystal violet

Observing bacteria under microscope

Observation of stained spores

Plantain roots staining to observe the association with arbuscular fungi

Isolation of spores in soil samples to check their viability

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations:

Around 15 Zip ties were used to repair the tarps that had been damaged by the wind.

Summary of any observatory issues: No observation today due to bad weather.

Summary of health and safety issues: Everyone was in good health today, with no issues reported. There were no safety concerns to note.

Questions, concerns and requests to Mission Support: None

Supplemental Operations Report -April 19th

Date: 4/19/2025
Name of person filing report: David Steinhour
Reason for Report: Routine
Non-Nominal Systems: Crew car. Power system battery, inverters, generator. Robotic observatory. HAB outer shell. Spirit.

Power system: "Solar: The battery bank does not hold charge when sun is down and low on the horizon. Inverter Slave 1 and 2 went offline and do not restart, which limits us to 4kW when on solar.
Main generator has been monitored for oil leaks; minor leaks observed, Moreover, the generator is consuming oil because of worn piston rings. Adding oil every day is necessary. Generator is limited to 8kW, see previous reports for details. Engine governor has been replaced, and the threads of one of the governor bolts was stripped and repaired with a heli-coil insert. Oil is still leaking from the flange that is held down by this fastener – the gasket could need replacing.
Main generator:
1) Oil, oil filter and air filter changed on 04/11/2025. Hours: 9260.2
2) Current hours – 9364.1"

Propane Readings: "Station Tank: 70%
Director Tank: 62%
Intern Tank: 84%
Generator Tank: 42%"

Water: "Hab Static Tank – 550 gallons
GreenHab – 200 gallons
Outpost tank – 225 gallons"

Rovers: "Sojourner rover used: Yes
Hours: 212.9
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: No
Notes on Rovers: Spirit is not charging. Perseverance’s parking brake needs to be adjusted."

Cars: "Hab Car used and why, where: To Hanksville for supplies.
Crew Car used and why, where: By crew from Grand Junction.
General notes and comments: Crew car driver’s side front ball joint is bad. Low oil pressure is most likely caused by a bad sensor (it does not change due to RPM or temperature like it should)."

Summary of Internet: Nominal
EVA suits and radios: "Suits: Nominal.
Comms: Nominal.
T-Echo EVA-link: Nominal"
Campus wide inspection, if action taken, what and why: Nominal
Summary of Hab Operations: Small cracks in the wall of the Hab in the loft area are allowing cold air intrusion. Static tank and pipe heaters unplugged. Thermostat setup for 60F due to warm weather.
Summary of GreenHab Operations: Nominal
Summary of SciDome Operations: Dual split is being turned on during day time for cooling.
Summary of Observatories Operations: Robotic observatory offline.
Summary of RAM Operations: Nominal
Summary of Outpost Operations: Static tank heater unplugged and removed. Screen doors installed in both trailers. Intern trailer heater turned off due to warm weather.
Summary of Health and Safety Issues: Nominal

Copyright © The Mars Society. All rights reserved. | Main Site